Find the equivalent fraction of 36 / 48 with
denominator 4

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Given, denominator = 4
∴N / 4 = 36 / 48
N × 48 = 4 × 36
N = (4 × 36) / 48
=3
Hence, the equivalent fraction is 3 / 4

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Let the numerator of equivalent fraction be x,
 
X/4 = 36/48
X/4 = 3/4
 
Therefore, X = 3
 
Hope it helps!
Thank you